# Quantitative assessment of biological dynamics with aggregate data

### v1

@@ 2.3 Data acquisition
The in vivo data shown in Results are obtained with the following methodology. Axenic cultures of Prochlorococcus strain MIT9215 cyanobacteria were maintained in an artificial seawater medium, AMP-MN [11], a derivative of AMP-A medium [44, 43] without any nitrogen amendment. Purity tests to determine the axenicity of cyanobacteria stocks and experimental cultures were performed routinely as previously described in [45]. All experiments were carried out at 24C in Percival I36VLX incubators (Percival, Boone, IA) with modified controllers that allowed for gradual increase and decrease of cool white light to simulate sunrise and sunset with peak midday light intensity of 150 $\mu$mol quanta m-2s-1 on a 14 hr:10 hr light:dark cycle [69]. The abundance of cyanobacteria was quantified by flow cytometry using a Guava EasyCyte 8HT flow cytometer (Millipore, Burlington, MA) with populations of Prochlorococcus determined by their red fluorescence [45, 14]. Raw measurements were obtained in batches of size $K=24$ for a total of $N=9$ time points spread over the duration of the experiment.
@@ 3 Results
In this section, we show how our modeling framework performs on estimating the parameters of IVPs. To demonstrate its effectiveness in revealing the correct parameter values, we first validate our model on synthetic data mimicking real ones that are generated with prescribed parameter values. Our in silico data are obtained by simulating the model of section 2.2. Subsequently, we demonstrate that our methods maintain their performance on real data. Our in vitro data are acquired as described in section 2.3. We also compare against naive parameter estimation methods based on least-squares fitting.
@@ 3.1 In silico growth curve data
To generate the synthetic data, we employ the IVP in eqs. 10 and 9 modeling Pro growth with the ground-truth parameter values $(P,Q,m,a)=(13,13\times 10^{4},2,55\times 10^{-6})$ which we chose in agreement with the in vitro data of the next section. Then, we generate cell density measurements $y_{1:N}^{1:K}$ according to eq. 12 and derive summary statistics $z_{1:N}^{1:M}$ calculated by eqs. 2 and 3. Our data are shown in fig. 2 with the upper panels corresponding to a scenario in which only batch means are maintained and the lower panels to a scenario where both batch means and variances are maintained after aggregation. We then employ our statistical learning framework to generate samples of the posterior distribution considering only sample means $p(P,Q,m,a|z_{1:N}^{1})$ or sample means and variances $p(P,Q,m,a|z_{1:N}^{1:2})$.
@@
Figure 2: Fitting the synthetic data with our statistical learning framework. The first row shows our results for a linear (mean only) constraint, and the second row shows our results for a quadratic (mean and variance) constraint. (top and bottom left) We use our MCMC sampled parameter values to generate thousands of ODE solutions which are plotted as the blue curves, we then use the MAP estimate to generate the red curve and all this is compared to the data points shown as black squares. (top and bottom right) while doing the parameter estimation we can also see how well we reproduce the hidden measurements. We show scatter plots of each parameter compared with the other in the off-diagonal panels. On the diagonal panels we show the histogram created by thousands of MCMC samples of each parameter and overlay via the red vertical line the ground truth parameter value that was used to generate the synthetic data. Here our results show good performance recovering the parameter values and therefore producing a well-fitting ODE solution.
To allow for comparison with ground truth, for each case, we approximate the maximum a posterior estimate (MAP) of our parameters by selecting the MCMC sample with the highest posterior probability density calculated by eq. 18, [62]. These estimates represent our framework’s best choices for parameter values in each case and are indicated by solid lines in fig. 2. Their specific values are $(\hat{P},\hat{Q},\hat{m},\hat{a})=(13.58,13.08\times 10^{4},2.21,54.80\times 10^{-6})$ for the first case considering only batch means and $(\hat{P},\hat{Q},\hat{m},\hat{a})=(13.33,12.89\times 10^{4},1.89,54.73\times 10^{-6})$ for the second case considering both batch means and variances. Both estimators are in good agreement with the ground truth.
@@
For the two panels in the left column of fig. 2, we also show a collection of sample trajectories. These are randomly selected MCMC samples that correspond to the solutions of eqs. 9 and 10. For both cases, we get a spread of trajectories around the MAP solution that indicates uncertainty coming from the assumed noise in the measurements and missing information due to data aggregation. The uncertainty in the upper left panel is higher than that in the lower left panel, as indicated by the wider spread of the sampler trajectories around MAP. This is expected behavior because we assimilate the same information (i.e. batch means) plus additional information (i.e. batch variances) in the second case.
@@
The two panels in the right column of fig. 2 show MCMC samples of the parameters. Specifically, along the diagonals, we show the MCMC approximations (histograms) of the marginal posteriors $p(Q|z_{1:N}^{1:M}),p(P|z_{1:N}^{1:M}),\allowdisplaybreaks\allowbreak p(m|z_{1:N}^{1:M}),p(a|z_{1:N}^{1:M})$ along with the MAP estimates as vertical lines. Again, although there is general agreement with the ground truth, there is more uncertainty in the case of only the batch mean than in the case of both batch mean and variance, as indicated by wider histograms. In the off-diagonal panels, we show MCMC approximations (scatter plots) of each pairing of the sample parameters. In both cases, our framework reveals preferences between combinations of parameters, indicating that, to comply with the supplied data, trajectories of the underlying IVP require specific values. For instance, the pair $(Q,a)$ has the highest correlation among all pairs. This is expected behavior, as our dynamical model, see eqs. 9 and 10, exhibits a degree of nonidentifiability between the initial amount of nutrients and nutrient affinity.
@@
In addition to showing that our framework successfully recovers the values of the ground truth parameters with quantified uncertainty, in fig. 3 we demonstrate the strength of our methods when challenged with scenarios involving more demanding data. To this end, we apply our learning framework on the same two cases (only batch mean and both batch mean and variance) by now considering scenarios of decreasing batch size $K=24,12,6,3$. In this way, we simulate a series of aggregate data generated in experiments of successively fewer raw measurements. Although fewer raw measurements result in increasingly noisier aggregate data, as seen in fig. 3 (upper two rows), our framework’s estimates remain in good agreement with the ground truth.
@@
A head-to-head comparison with naive parameter estimation procedures mediated by least-square fitting (LS), see appendix D, indicates superior performance. In particular, in fig. 3 (bottom two rows) and table 1 we quantify the error percentage in our MAP estimates and those resulting from LS for each parameter $X=Q,P,m,a$ using
@@
%$\%\text{Error}=\left(\frac{X^{\text{\rm estimate}}}{X^{\text{\rm ground}}}-1\right)\times 100\%.$
@@
While the error for all three methods increases as batch size $K$ decreases, and so the noise persisting in the aggregated data increase, LS consistently produces the least accurate estimates. This indicates that, in addition to recovering the ground truth more accurately, our learning framework is also more robust when faced with excessively noisy data.
@@
Figure 3: Batch size robustness of our methods. We show the effect of reducing the batch size $K$ from $24$ to $3$ going from left to right across all figures. In the upper two rows, we show the solutions of our IVP eqs. 9 and 10 for the linear constraints of the batch mean (top) and the quadratic constraint of the batch mean and variance (bottom) as $K$ is decreased. In the bottom two rows, we show the percentage error between the linear constraint MAP predictions (blue), quadratic constraint MAP prediction (red), and a LS prediction (yellow) for the parameter values $(P,Q,m,a)$ compared to their ground truth. This is shown in the upper panels as their individual errors in the form of a bar chart and in the bottom panel as a line plot of the total error (sum of individual errors) for each batch size. As we reduce the number of samples per time point down from a collection of experiments $K=24$, to one experiment $K=3$ we see the least squares method break down whereas our method is better at handling the decrease in the signal to noise. See table 1 in appendix D for a table of the values used in the least squares comparison.
Units Ground truth Only batch means Both batch means and variances Least squares
@@
$P$ $\text{cells}/m^{3}$ $13\times 10^{4}$ $12.95\times 10^{4}$ 0.42 $13.43\times 10^{4}$ 3.34 $10.43\times 10^{4}$ 19.75
$Q$ $1/\text{days}$ 13 13.15 1.15 12.77 1.81 0.021 99.84
$m$ $1/\text{days}$ 2 1.83 8.33 2.02 1.13 0.308 84.6
$a$ $m^{3}/(\text{cells}\cdot\text{days})$ 5.5 ×$\times 10^{-6}$ 5.51 ×$\times 10^{-6}$ 0.24 5.33 ×$\times 10^{-5}$ 3.00 5.10 ×$\times 10^{-6}$ 189.10
@@
$P$ $\text{cells}/m^{3}$ $13\times 10^{4}$ $13.52\times 10^{4}$ 3.97 $13.52\times 10^{4}$ 0.633 $11.92\times 10^{4}$ 8.28
$Q$ $1/\text{days}$ 13 12.45 4.23 12.45 7.08 4.25 67.31
$m$ $1/\text{days}$ 2 1.81 9.73 1.81 2.52 4.20 110.0
$a$ $m^{3}/(\text{cells}\cdot\text{days})$ $5.5\times 10^{-6}$ $5.27\times 10^{-6}$ 4.26 $5.27\times 10^{-6}$ 0.954 $6.50\times 10^{-6}$ 18.18
@@
$P$ $\text{cells}/m^{3}$ $13\times 10^{4}$ $12.32\times 10^{4}$ 5.22 $13.77\times 10^{4}$ 5.96 $12.48\times 10^{4}$ 4.02
$Q$ $1/\text{days}$ 13 13.08 0.627 13.02 0.124 4.59 64.70
$m$ $1/\text{days}$ 2 1.87 5.68 2.00 0.193 8.93 17.45
$a$ $m^{3}/(\text{cells}\cdot\text{days})$ $5.5\times 10^{-6}$ $5.69\times 10^{-6}$ 3.40 $5.22\times 10^{-6}$ 5.07 $6.46\times 10^{-6}$ 17.45
@@
$P$ $\text{cells}/m^{3}$ $13\times 10^{4}$ $11.80\times 10^{4}$ 9.21 $12.86\times 10^{4}$ 1.11 $14.30\times 10^{4}$ 10.03
$Q$ $1/\text{days}$ 13 30.56 135.11 34.46 165.07 63.84 391.08
$m$ $1/\text{days}$ 2 2.008 0.424 2.01 0.620 18.0 800.0
$a$ $m^{3}/(\text{cells}\cdot\text{days})$ $5.5\times 10^{-6}$ $5.94\times 10^{-6}$ 8.09 $5.75\times 10^{-6}$ 4.58 $3.90\times 10^{-6}$ 29.09
Table 1: Quantitative validation against ground truth. The comparison shows that our model has better agreement with the ground truth than least squares, particularly as batch size $K$ size decreases.
3.2 In vitro growth curve data
Having demonstrated our framework’s ability to accurately recover ground-truth parameter values with synthetic data, we now demonstrate its application on real laboratory data. Here, our summary statistics $z_{1:N}^{1:M}$ are obtained directly from the batch growth Pro data. Our data are shown in fig. 4 and, as previously, we distinguish a case that considers only batch means (upper panels) and one that considers both batch means and variances (lower panels). For the two cases, our MAP estimators are $(\hat{P},\hat{Q},\hat{m},\hat{a})=(30.89,12.45\times 10^{4},1.94,57.43\times 10^{-6})$ and $(\hat{P},\hat{Q},\hat{m},\hat{a})=(53.16,12.13\times 10^{4},1.85,55.51\times 10^{-6})$, respectively. These represent growth rate estimates in line with the empirical literature (c.f. Figure 2 of [41], Figure S5 of [27]). In addition to MAP estimates, our framework fully quantifies uncertainty in this case in either trajectories or parameter values. As anticipated, our learning framework also performs well with real Pro experimental data.
Figure 4: Fitting real data from Pro experiments. The first row shows our results for a linear constraint and the second row shows our results for a quadratic constraint. (top and bottom left) We use our MCMC sampled parameter values to generate thousands of ODE solutions which are plotted as the blue curves, we then use the MAP estimate to generate the red curve and all this is compared to the data points shown as black squares. (top and bottom right) while doing the parameter estimation we can also see how well we reproduce the hidden measurements. We show scatter plots of each parameter compared with the other in the off-diagonal panels. On the diagonal panels we show the histogram created by thousands of MCMC samples of each parameter.
@@ 3.3 Sensitivity analysis
Our Bayesian framework depends on prior probability distributions, which may influence our posterior estimates. To test the effect of our prior’s influence on the resulting estimates, we conduct a sensitivity analysis taking into account changes in the hyperparameters of eqs. 13, 14, 15 and 16. Specifically, we vary the expected values of the prior represented by $\psi_{Q},\psi_{P},\psi_{m},\psi_{a}$ within ±$\pm 10\%$ of their baseline and obtain the resulting marginal posteriors $p(Q|z_{1:N}^{1:M}),p(P|z_{1:N}^{1:M}),p(m|z_{1:N}^{1:M}),p(a|z_{1:N}^{1:M})$ over the individual parameters. As shown on fig. 5 (upper panels), hyperparameter changes are not transmitted to the resulting posteriors, indicating that parameter point estimates are robust with respect to the model’s hyperparameters. In addition, a quantitative comparison of the estimated trajectories $p^{\hat{g}}(\cdot)$ via the relative root-mean-square error
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{n}\left(\frac{p^{g}(t_{n})}{z_{n}}-1\right)^{2}},$
also shown on fig. 5 (lower panel), illustrates that hyperparameter changes are not transmitted to the trajectories as well, indicating that our recovered dynamics are informed by the supplied data rather than the hyperparameter choices.
Figure 5: Sensitivity analysis of our model’s hyper-parameters for $Q,P,m,a$. To evaluate the sensitivity, we change the values of the hyperparameter $\psi_{Q},\psi_{P},\psi_{m},\psi_{a}$ from a baseline (blue) to an increase (red) of $10\%$ and a decrease (green) of $10\%$. In the upper panels, we show marginal posteriors for each parameter across the three cases. In the lower panel, we show the root mean squared error (RMSE) when we compare our MCMC trajectories from the three different cases with the data. Across all the figures, we see a significant overlap of the three cases, indicating that our methods are not sensitive to changes in the hyper-parameter values.

### v2

@@ 2.3 Data acquisition
The in vivo data shown in Results are obtained with the following methodology. Axenic cultures of Prochlorococcus strain MIT9215 cyanobacteria were maintained in an artificial seawater medium, AMP-MN [12], a derivative of AMP-A medium [54, 53] without any nitrogen amendment. Purity tests to determine the axenicity of cyanobacteria stocks and experimental cultures were performed routinely as previously described in [55]. All experiments were carried out at 24C in Percival I36VLX incubators (Percival, Boone, IA) with modified controllers that allowed a gradual increase and decrease of cool white light to simulate sunrise and sunset with a peak midday light intensity of 150 $\mu$mol quanta m-2s-1 on a 14 hr:10 hr light:dark cycle [91]. The abundance of cyanobacteria was quantified by flow cytometry using a Guava EasyCyte 8HT flow cytometer (Millipore, Burlington, MA) with populations of Prochlorococcus determined by their red fluorescence [55, 15]. Raw measurements were obtained in batches of size $K=24$ for a total of $N=9$ time points spread over the duration of the experiment.
@@ 3 Results
In this section, we show how our framework performs on estimating the parameters of IVPs. To demonstrate its effectiveness in revealing the correct parameter values, we first validate our model on synthetic data, mimicking the characteristics of real data, that are generated with prescribed parameter values. Our in silico experiments are conducted by simulating the model of section 2.2. Subsequently, we demonstrate that our methods maintain their performance on real laboratory data. Our in vivo experiments are conducted as described in section 2.3. We also compare against naive parameter estimation methods based on least-squares fitting.
@@ 3.1 In silico growth curve data
[…]
@@
Figure 2: Fitting in silico data with our statistical learning framework. The first row shows our results for a mean only constraint, and the second row shows our results for a mean and standard deviation constraint. Left: We use our MCMC parameter values to generate IVP solutions which are plotted as the blue curves, we then use the MAP estimate to generate the red curve and compare to the data points shown in black. Right: in the off-diagonal panels, we show MCMC scatter plots of pairwise parameter comparisons and, in the diagonal, we show MCMC histograms of each parameter and overlay the ground truth values with vertical red lines. Our results recover well the parameter truth parameters values and producing well-fitting IVP solutions.
[…]
@@
For the two groups of panels in the left column of fig. 2, we also show a collection of sample trajectories. These are randomly selected MCMC samples that correspond to the solutions of eqs. 9 and 10. For both cases, we obtain a spread of trajectories around the MAP solution that indicates uncertainty coming from the noise in the measurements and missing information due to data aggregation. The uncertainty in the upper left panel is higher than that in the lower left panel, as indicated by the wider spread of the sampler trajectories around MAP. This is expected behavior because we assimilate the same information (i.e. batch means) plus additional information (i.e. batch standard deviations) in the second case.
@@
The right panels in fig. 2 show MCMC samples of the parameters. Specifically, along the diagonals, we show the MCMC approximations (histograms) of the marginal posteriors $\mathcal{P}(Q|z_{1:N}^{1:M}),\mathcal{P}(P|z_{1:N}^{1:M}),\allowbreak\mathcal{P}(m|z_{1:N}^{1:M}),\mathcal{P}(a|z_{1:N}^{1:M})$ along with the MAP estimates as vertical lines. Again, although there is general agreement with the ground truth, there is more uncertainty in the case of only the batch means than in the case of both batch means and standard deviations, as indicated by wider histograms. In the off-diagonal panels, we show MCMC approximations (scatter plots) of each pairing of the sample parameters. In both cases, our framework reveals preferences between combinations of parameters, indicating that, to comply with the supplied data, the trajectories of the underlying IVP require specific configurations of parameter values. For instance, the pair $(m,a)$ has the highest correlation among all pairs. This is expected behavior, as our dynamical model, see eqs. 9 and 10, depends only on the ratio $m/a$ and not separately on $a$; therefore, the value of $a$ can be estimated only relative to the value of $m$, as seen.
@@
In addition to showing that our framework successfully recovers the values of the ground truth parameters with quantified uncertainty, in fig. 3 we demonstrate the performance of our methods when challenged with scenarios involving more demanding data. To this end, we apply our learning framework in the same two cases (only batch means and both batch means and standard deviations) now considering scenarios of decreasing batch size $K=24,12,6,3$. In this way, we simulate a series of aggregate data generated in experiments of successively fewer raw measurements. Although fewer raw measurements result in increasingly noisier aggregate data, as seen in fig. 3 (upper two rows), our framework’s estimates remain in good agreement with the ground truth. This indicates our framework’s robustness to increased noise.
@@
A head-to-head comparison with naive parameter estimation procedures mediated by least-square fitting (LS), see appendix D, indicates superior performance. In particular, in fig. 3 (bottom two rows) and table 1 we quantify the percentage error in our MAP estimates and those resulting from LS for each parameter. For clarity, our error metrics are given by
@@
%$\%\text{Error}=\left|\frac{X^{\text{\rm estimate}}}{X^{\text{\rm ground}}}-1\right|\times 100\%,$
@@
where $X$ stands for any of $Q,P,m,a$. As seen, while the error for all three methods generally increases as the batch size $K$ decreases, and therefore the noise that persists in the aggregated data increases, LS consistently produces the least accurate estimates. In contrast, our framework consistently produces the most accurate ones. This indicates that, in addition to recovering the ground truth more accurately, our learning framework is also more robust when faced with excessively noisy data than existing practices.
Furthermore, a comparison between the batch means only scenario and the scenario incorporating both batch means and standard deviations, afforded by our framework, indicates that the best estimates are consistently obtained with the latter. Again, this is not surprising, since the latter scenario assimilates the largest amount of data.
@@
Figure 3: Batch size robustness of our methods. We show the effect of reducing the batch size $K$ from $24$ to $3$ going from left to right across all figures. In the upper two rows, we show the solutions of our IVP eqs. 9 and 10 for the constraints on the batch means (top) and the constraints on the batch means and standard deviations (bottom) as $K$ is decreased. In the bottom two rows, we show the percentage error between the mean only constraint MAP predictions (blue), mean and standard deviation constraint MAP prediction (red), and a LS prediction (yellow) for the parameter values $(Q,P,m,a)$ compared to their ground truth. This is shown in the upper panels as their individual errors in the form of a bar chart and in the bottom panel as a line plot of the total error (sum of individual errors) for each batch size. As we reduce the number of samples per time point down from a collection of experiments $K=24$, to one experiment $K=3$ we see LS breaks down whereas our framework copes better with the decrease in the signal to noise content of the resulting data. See table 1 for the precise values used in the comparison.
Ground Only batch means Both batch means and Least squares
Units truth standard deviations
@@
$m$ $1/\text{days}$ 0.5 0.446 10.828 0.478 4.300 0.529 5.811
$a$ $\text{ml}/(\text{cells}\cdot\text{days})$ 0.00001 .0000102 2.187 .000010078 0.780 .00000896 10.372
@@
[…]
$m$ $1/\text{days}$ 0.5 0.405 19.074 0.423 15.462 0.335 33.054
$a$ $\text{ml}/(\text{cells}\cdot\text{days})$ 0.00001 .0000101 1.838 .0000100 0.198 .0000103 3.898
[…]
$m$ $1/\text{days}$ 0.5 0.502 0.313 0.491 1.750 0.569 13.892
$a$ $\text{ml}/(\text{cells}\cdot\text{days})$ 0.00001 .00000955 4.406 .00000966 $3$0000115 15.797
[…]
$m$ $1/\text{days}$ 0.5 0.354 29.120 0.481 3.862 3.475 595.081
$a$ $\text{ml}/(\text{cells}\cdot\text{days})$ 0.00001 .00000788 21.138 .00000795 20.484 .0000115 15.305
Table 1: Quantitative validation against ground truth. The comparison shows that our model outperforms LS estimation, particularly as batch size $K$ size decreases.
3.2 In vivo growth curve data
[…]
Figure 4: Fitting in vivo data with our statistical learning framework. The first row shows our results for a mean only constraint, and the second row shows our results for a mean and standard deviation constraint. Left: We use our MCMC parameter values to generate IVP solutions which are plotted as the blue curves, we then use the MAP estimate to generate the red curve and compare to the data points shown in black. Right: in the off-diagonal panels, we show MCMC scatter plots of pairwise parameter comparisons and, in the diagonal, we show MCMC histograms of each parameter. Our results generate parameter values in line with expectations based on the literature. We also note the ridge like structure of the $(m,a)$ joint distribution indicating a non-identifiability as moves along one domain are highly correlated with the other to maintain a fixed value for $\frac{m}{a}$ in the IVP.
@@ 3.3 Sensitivity analysis
Because our Bayesian framework depends on prior probability distributions, which may influence our posterior estimates, for all analyses shown we choose weakly informative priors for the parameters [21]. We achieve these using in all simulations default values for the shape hyperparameters $\phi_{Q}=\phi_{P}=\phi_{m}=\phi_{a}=2$ in eqs. 13, 14, 15 and 16. However, the expectation hyperparameters $\psi_{Q},\psi_{P},\psi_{m},\psi_{a}$ may still have a substantial effect. For this reason, to quantify our prior’s influence on the resulting estimates, we conduct a sensitivity analysis taking into account large changes in their values.
Specifically, we vary the values of $\psi_{Q},\psi_{P},\psi_{m},\psi_{a}$ within ±$\pm 50\%$ of their baseline values used in our earlier simulations and derive the resulting marginal posteriors $\mathcal{P}(Q|z_{1:N}^{1:M}),\mathcal{P}(P|z_{1:N}^{1:M}),\mathcal{P}(m|z_{1:N}^{1:M}),\mathcal{P}(a|z_{1:N}^{1:M})$. As shown in fig. 5 (top panels), our hyperparameter changes are not transmitted to the resulting posteriors, indicating that the associated point estimates are robust, as expected, to the choice of the priors.
In addition, a quantitative comparison of the estimated MAP trajectories $p^{\hat{g}}(\cdot)$ via the relative root-mean-square error
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{n}\left(\frac{p^{\hat{g}}(t_{n})}{z_{n}}-1\right)^{2}},$
also shown on fig. 5 (bottom panel), illustrates that hyperparameter changes are not transmitted to the trajectories as well. This indicates that the dynamics recovered by our framework are informed by the supplied data rather than by the hyperparameters.
Figure 5: Sensitivity analysis of our posterior estimates of the hyperparameters. To evaluate the sensitivity, we change the values of the hyperparameter $\psi_{Q},\psi_{P},\psi_{m},\psi_{a}$ from a baseline (blue) to an increase (red) of $50\%$ and a decrease (green) of $50\%$. In the upper panels, we show marginal posteriors for each parameter across the three cases. In the lower panel, we show the root mean squared error (RMSE) when we compare our MCMC trajectories from the three different cases with the data. Across all panels, we see a significant overlap of the three cases, indicating that our methods are insensitive to changes in the hyperparameter values.